# Virne: A Comprehensive Benchmark for Deep RL-based Network Resource Allocation in NFV

### v1

@@ C.3.2 Implemented Traditional Algorithms
MIP [5] Mixed-Integer Programming (MIP)
R-Rounding [67] Random Rounding
D-Rounding [67] Deterministic Rounding
RW-Rank [68] Random Walk (RW)
GRC-Rank [7] Global Resource Control (GRC)
NRM-Rank [8] Node Resource Management (NRM)
NEA-Rank [9] Node Essentiality Assessment (NEA)
PL-Rank [51] Priority of Location (RL)
GA-Meta [55] Genetic Algorithm (GA)
PSO-Meta [57] Particle Swarm Optimization (PSO)
ACO-Meta [54] Ant Colony Optimization (ACO)
SA-Meta [69] Simulated Annealing (SA)
TS-Meta [70] Tabu Search (TS)
@@
  • MIP [5]: Mixed-Integer Programming (MIP) is an exact method that provides optimal solutions by solving a system of linear equations with integer constraints.
  • R-Rounding [67]: A rounding method that applies random rounding to generate approximate solutions for NFV-RA.
  • D-Rounding [67]: This method uses deterministic rounding to map continuous variables to discrete values while providing an approximation to the optimal solution.
@@
  • NRM-Rank [8]: A heuristic method based on Node Resource Management (NRM) metric. This approach ranks virtual and physical nodes before mapping with a greedy method.
  • GRC-Rank [7]: This method uses a Global Resource Control (GRC) strategy based on random walks for node ranking. After ranking, the nodes are mapped accordingly.
  • NEA-Rank [9]: It ranks nodes using the Node Essentiality Assessment (NEA) metric.
  • RW-Rank [68]: A node ranking approach based on random walks to estimate node priorities in the embedding process.
  • PL-Rank [51]: This algorithm considers node proximity and evaluates paths comprehensively, ranking nodes based on their location priorities and the overall efficiency of the physical infrastructure.
@@
  • GA-Meta [55]: A meta-heuristic method based on genetic algorithms. It models each node mapping solution as a chromosome and iteratively explores the solution space by simulating the process of genetic selection and evolution.
  • PSO-Meta [56]: A meta-heuristic using Particle Swarm Optimization (PSO), where particles explore the NFV-RA solution space by adjusting their positions based on their own and neighbors’ experiences.
  • ACO-Meta [54]: A meta-heuristic based on Ant Colony Optimization (ACO) that simulates ant foraging behavior to explore the NFV-RA solution space and update pheromone trails to guide the search.
  • SA-Meta [69]: A meta-heuristic using Simulated Annealing (SA) to approximate the global optimum, exploring the NFV-RA solution space with both upward and downward transitions to escape local minima.
  • TS-Meta [70]: A meta-heuristic based on Tabu Search (TS), which uses a memory structure to store visited solutions, promoting diversity and avoiding local optima.
C.4 Evaluation Metric Definitions
@@
    where $\mathcal{T}$ denotes the total operational time of the network system. $\mathcal{I}(t)$ is the set of all VN requests arriving at time slot $t$, and ~$\tilde{\mathcal{I}}(t)$ is the subset of those requests that are accepted. The notation ||$\cdot$|| ⋅ | denotes the cardinality of a set.
@@ D.1.1 Implementation Settings
Algorithm Implementation. The implementation of the RL-based NFV-RA algorithm is based on PyTorch [71]. Regarding neural policies (e.g., MLP, CNN, GNN, etc), we generally set their layers to 3 and the hidden dimension of neural networks to 128. For RL optimization, we set the reward discount factor $\lambda$ to 0.99 and use the Adam optimizer with a learning rate of 0.001 and a batch size of 128. The action section during training and testing follows a sampling strategy and a greedy strategy, respectively. Additionally, we employ the $k$-shortest paths algorithm with $k=10$ for link routing, selecting the shortest physical path that meets the bandwidth requirements for each virtual link.
Training and Testing. For RL-based methods, we train policies for each average arrival rate $\eta$, with random initialization of seeds in every simulation. The number of training epochs is set to 50, which is sufficient for all algorithms to converge. Typically, training a deep RL-based NFV-RA method completes within 5 hours. During testing, we evaluate the performance of all algorithms by repeating each test with 10 different random seeds (i.e., 0, $1$, $2$, $\ldots$, $9$) for each average arrival rate $\eta$, thereby ensuring statistical significance.
@@ D.1.2 Evaluated Physical Topologies
WX100 is a medium-scale network generated by the Waxman method [36]. It consists of 100 nodes connected by approximately 500 links, resulting in a density of 0.05.
@@ D.2.1 Evaluation on Varying Traffic Rates
As illustrated in Figure 6 (a), (d), and (g), there is a general downward trend in the Request Acceptance Rate (RAC) for all algorithms as the average arrival rate $\eta$ increases across all three topologies. This is an expected outcome, as higher traffic intensity leads to increased competition for finite physical network resources, inevitably resulting in more VN request rejections. For instance, on WX100, the RAC for PPO-DualGAT decreases from 0.99 at $\eta=0.04$ to 0.781 at $\eta=0.16$, and further to 0.66 at $\eta=0.28$. In comparison, MCTS drops notably.
Regarding the Long-term Revenue-to-Cost ratio (LRC), shown in Figure 6 (b), (e), and (h), we observe that 5 many RL-based algorithms, particularly PPO-DualGAT and PPO-DualGCN, demonstrate remarkable stability in LRC across varying traffic rates. For example, on WX100, PPO-DualGAT’s LRC remains very stable, i.e., 0.778 at $\eta=0.04$, 0.737 at $\eta=0.16$, and 0.746 at $\eta=0.28$. This indicates that while they accept fewer requests under high load, the quality of accepted embeddings in terms of resource efficiency remains consistently high. In contrast, MCTS’s LRC on WX100 declines more significantly from 0.543 ($\eta=0.04$) to 0.444 ($\eta=0.16$) and 0.410 ($\eta=0.28$). Traditional heuristics like PL-Rank also show relatively stable LRC (e.g., on WX100, 0.662 at $\eta=0.04$, 0.668 at $\eta=0.16$, 0.650 at $\eta=0.28$) but generally at a lower level than top RL performers.
The results on Long-term Average Revenue (LAR) are depicted in Figure 6 (c), (f), and (I). Generally increasing with $\eta$, most algorithms process more VN requests. Algorithms like PPO-DualGAT and PPO-DualGCN consistently achieve higher LAR across the range of $\eta$ values, showcasing their superior capability in maximizing revenue even as conditions change.
These experiments highlight that 6 while some algorithms (often sophisticated RL policies like PPO-DualGAT) exhibit graceful performance degradation and maintain high efficiency (LRC), others are more sensitive to load changes. This allows researchers and practitioners to select algorithms that are well-suited to the anticipated traffic dynamics of their target environments. For instance, for networks expecting highly variable loads, algorithms demonstrating flatter RAC and LRC curves across $\eta$ would be preferable. The consistent outperformance of dual GNN architectures (PPO-DualGAT and PPO-DualGCN) further corroborates the findings in the main paper regarding their effectiveness.
@@ D.2.2 Evaluation on Fluctuating Demand Distribution
As illustrated in Figure 7 (a), (d), and (g) for RAC, most algorithms exhibit an initial sharp decrease, particularly as the simulation transitions through phases with increasing VN resource demands. This phase quickly consumes available physical resources, leading to lower acceptance rates. As the simulation progresses into phases where VN sizes increase, the RAC for many algorithms continues to decline or stabilizes at a lower level. This is because larger and more complex VNs are inherently more challenging to embed. Notably, 7 advanced RL methods like PPO-DualGAT and PPO-DualGCN tend to maintain a higher RAC compared to simpler MLP-based policies or traditional heuristics throughout these fluctuations, indicating better adaptation.
The LRC metric, shown in Figure 7 (b), (e), and (h), also reflects the algorithms’ adaptability. Many algorithms show a dip in LRC when the demand characteristics change, but some, like PPO-DualGAT and PPO-DualGCN, manage to sustain a relatively higher LRC. For WX100 ($\eta=0.04$) under changeable demands, PPO-DualGAT records an LRC of 0.703 and PPO-DualGCN an LRC of 0.700, whereas PPO-MLP has an LRC of 0.643. This suggests they are better at finding resource-efficient embeddings even when faced with unfamiliar or more challenging VN request types.
The Total Revenue (i.e., cumulative LAR), depicted in Figure 7 (c), (f), and (i), shows a clear differentiation in the algorithms’ ability to accumulate revenue under these dynamic conditions. Algorithms that adapt well and maintain higher RAC and LRC will naturally accrue more total revenue. PPO-DualGAT and PPO-DualGCN consistently demonstrate a steeper accumulation of total revenue. For WX100 ($\eta=0.04$) under changeable demands, PPO-DualGAT achieves a long-term average time revenue of 5080.33, compared to 4510.80 for PPO-MLP. This underscores their robustness and effectiveness in dynamic environments.
@@
These results demonstrate that policies trained on one specific distribution of VNs may not always generalize perfectly to others. This highlights 8 the importance of generalization not just to varying load, but also to varying types of demand. For practical systems where service characteristics can evolve, choosing algorithms that demonstrate robustness in such fluctuating scenarios, as identifiable through Virne, is important.
D.3 Solvability via Offline Evaluation
Traditional online testing of NFV-RA algorithms, while reflecting real-world scenarios, hinders direct comparisons of intrinsic algorithmic solvability due to the continuously evolving state of the PN and dynamic VN arrivals and departures. It becomes challenging to discern whether failures or poor performance stem from an algorithm’s inherent limitations or from transient, potentially unsolvable, network states. To address this, Virne provides a controlled offline evaluation environment. This environment utilizes a series of static instances, each comprising a PN and a VN, to assess an algorithm’s ability to find high-quality, feasible solutions under reproducible conditions. This section focuses on evaluating the solution quality, specifically the Revenue-to-Cost (R2C) ratio, achieved by different algorithms for VNs of varying sizes. The evaluation is conducted on a representative PN, WX100. The results are visualized in Figure 8, which presents a heatmap of average solution quality (R2C ratio) for VN sizes ranging from 2 to 10 nodes.
As shown in Figure 8, there is a general trend of decreasing solution quality as the size of the VN increases. This is expected, as larger VNs typically have more complex topologies and higher resource demands, making them more challenging to embed efficiently. Additionally, we can observe that 8 RL-based methods with advanced graph representations, particularly PPO-DualGAT and PPO-DualGCN, demonstrate superior solvability across most VN sizes. For instance, PPO-DualGAT consistently achieves high R2C ratios. These algorithms are often highlighted in red (best-performing) in Figure 8 for larger VN sizes. Additionally, 9 Simpler RL policies like PPO-MLP also show good performance, especially for smaller VNs, but degrade more noticeably for larger VNs due to their limited representation ability. Meanwhile, traditional heuristic algorithms show mixed results. NEA-Rank performs exceptionally well for very small VNs, achieving an R2C of 1.0 for 2-node VNs. However, its performance, along with PL-Rank, tends to be lower than the top RL methods as VN size increases. Meta-heuristic algorithms like PSO-Meta tend to underperform in terms of solution quality in this offline setting, especially for larger VNs.
This study allows for a granular understanding of each algorithm’s fundamental problem-solving strength (solvability) concerning VN complexity, isolated from the dynamics of online environments. The results indicate that 10 while some algorithms excel with smaller VNs, their performance can degrade significantly as complexity increases. This detailed solvability analysis helps researchers identify the strengths and weaknesses of different algorithmic approaches for varying scales of the problem. For instance, the superior performance of dual GNN architectures across different VN sizes suggests their strong capability in capturing complex relationships for effective embedding. Practitioners can leverage these insights to select algorithms that are best suited for the specific types and sizes of VNs they anticipate in their networks.
D.4 Scalability on Network Sizes

### v2

@@ C.3.2 Traditional Algorithm Implementation
MIP (Chowdhury et al., 2009) Mixed-Integer Programming (MIP)
R-Rounding (Chowdhury et al., 2012) Random Rounding
D-Rounding (Chowdhury et al., 2012) Deterministic Rounding
RW-Rank (Cheng et al., 2011) Random Walk (RW)
GRC-Rank (Gong et al., 2014) Global Resource Control (GRC)
NRM-Rank (Zhang et al., 2018) Node Resource Management (NRM)
NEA-Rank (Fan et al., 2023) Node Essentiality Assessment (NEA)
PL-Rank (Fan et al., 2021) Priority of Location (RL)
GA-Meta (Zhang et al., 2019) Genetic Algorithm (GA)
PSO-Meta (Jiang and Zhang, 2021) Particle Swarm Optimization (PSO)
ACO-Meta (Fajjari et al., 2011) Ant Colony Optimization (ACO)
SA-Meta (Wang et al., 2013) Simulated Annealing (SA)
TS-Meta (Wang et al., 2017) Tabu Search (TS)
@@
  • MIP (Chowdhury et al., 2009): Mixed-Integer Programming (MIP) is an exact method that provides optimal solutions by solving a system of linear equations with integer constraints.
  • R-Rounding (Chowdhury et al., 2012): A rounding method that applies random rounding to generate approximate solutions for NFV-RA.
  • D-Rounding (Chowdhury et al., 2012): This method uses deterministic rounding to map continuous variables to discrete values while providing an approximation to the optimal solution.
@@
  • NRM-Rank (Zhang et al., 2018): A heuristic method based on Node Resource Management (NRM) metric. This approach ranks virtual and physical nodes before mapping with a greedy selection method.
  • GRC-Rank (Gong et al., 2014): This method uses a Global Resource Control (GRC) strategy based on random walks for node ranking. After ranking, the nodes are mapped accordingly.
  • NEA-Rank (Fan et al., 2023): It ranks nodes using the Node Essentiality Assessment (NEA).
  • RW-Rank (Cheng et al., 2011): A node ranking approach based on random walks to estimate node priorities in the embedding process.
  • PL-Rank (Fan et al., 2021): This algorithm considers node proximity and evaluates paths comprehensively, ranking nodes based on their location priorities and the overall efficiency of the physical infrastructure.
@@
  • GA-Meta (Zhang et al., 2019): A meta-heuristic method based on genetic algorithms. It models each node mapping solution as a chromosome and iteratively explores the solution space by simulating the process of genetic selection and evolution.
  • PSO-Meta (Su et al., 2014): A meta-heuristic using Particle Swarm Optimization (PSO), where particles explore the NFV-RA solution space by adjusting their positions based on their own and neighbors’ experiences.
  • ACO-Meta (Fajjari et al., 2011): A meta-heuristic based on Ant Colony Optimization (ACO) that simulates ant foraging behavior to explore the NFV-RA solution space and update pheromone trails to guide the search.
  • SA-Meta (Wang et al., 2013): A meta-heuristic using Simulated Annealing (SA) to approximate the global optimum, exploring the NFV-RA solution space with both upward and downward transitions to escape local minima.
  • TS-Meta (Wang et al., 2017): A meta-heuristic based on Tabu Search (TS), which uses a memory structure to store visited solutions, promoting diversity and avoiding local optima.
C.4 Modular and Extensible Implementation Pipeline
To facilitate rapid implementation and community-driven development, Virne is with a highly modular and extensible implementation pipeline for RL-based methods. This design is guided by the principle of separation of concerns, where distinct functional components of the RL workflow can be independently developed, configured, and interchanged. This section outlines the architecture of this pipeline and a customized configuration example, as illustrated in Figures 7 and 8.
Figure 7: The modular architecture of the RL-based implementation pipeline in Virne. The pipeline is decoupled into three core, interchangeable modules: Instance Environment, RL Agent, and RL Trainer. Each module contains several sub-components (e.g., feature constructor, neural policy, learning algorithm) that can be individually customized and combined.
C.4.1 Modularity Description of Pipeline
As shown in Figure 7, our pipeline consists of the following high-level, interoperable modules.
Instance Environment simulates the NFV system and handles the agent’s interactions. It is designed with customizable components, including the Action Controller, which defines the step-wise logic for state transition in a gym-style environment; the Reward Calculator, which allows for different reward schemes to be easily plugged in.
Feature Constructor is responsible for processing the raw state information from the environment (PN and VN) and extracting meaningful features. It supports various feature engineering strategies, such as including basic node/edge attributes or augmenting them with advanced topological metrics.
Neural Policy represents the core of the agent’s learned strategy. Our pipeline supports a wide array of Neural Networks, from simple MLPs to more complex structures like CNNs, RNNs, and various GNNs. The user can also customize the Decoding Strategy that translates the probabilistic output of the neural policy into a concrete action.
RL Trainer manages the learning process, which includes the rollout buffer that stores the interaction experiences (state, action, reward, etc.) used for training; and Learning Algorithms: that is used to update the agent’s policy. Virne integrates a comprehensive suite of solvers, such as PG, A2C, PPO, and more advanced options like SafeRL and MetaRL solvers.
C.4.2 Extensibility in Practice
This modular design enables users to develop new algorithms flexibly, as shown in Figure 8.
Configuration-driven Customization: Users can easily assemble different pipelines by modifying configuration files. As shown in the top part of Figure 8, parameters like the reward calculator or the types of features to use can be set with simple string or boolean flags. This enables extensive experimentation without writing any new code.
Class-based Extension: For more advanced customization of flexible components, as shown in the bottom part of Figure 8, users can create custom environments or solvers (e.g., CustomEnv, CustomSolver) by inheriting from base classes. New components, such as a novel policy architecture or a custom RL solver, can be injected as callable functions or classes during instantiation. This powerful design pattern significantly lowers the barrier for implementing and testing new research ideas within a standardized framework.
Figure 8: A code example of implementing a new NFV-RA algorithm via module customization.
C.5 Evaluation Metric Definitions
@@
    where $\mathcal{T}$ denotes the total operational time of the network system. $\mathcal{I}(t)$ is the set of all VN requests arriving at time slot $t$, and $\tilde{\mathcal{I}}(t)$ is the subset of those requests that are accepted. The notation $|\cdot|$ denotes the cardinality of a set.
@@
C.6 Evaluation Protocol Definitions
To provide a rigorous basis for comparing algorithmic performance across diverse NFV scenarios, we formally define the three core evaluation perspectives as follows:
  • Solvability ($\mathcal{S}$): This metric quantifies an algorithm’s ability to satisfy the hard constraints of the NFV problem (e.g., CPU, bandwidth, and latency limits) in the offline scenario. It is formally defined as the acceptance ratio:
    $\mathcal{S}=\frac{N_{success}}{N_{total}}$ (14)
    where $N_{success}$ is the number of successfully embedded VNRs and $N_{total}$ is the total number of arrived requests.
  • Generalization ($\mathcal{G}$): We define generalization conceptually as the robustness of a policy $\pi_{\theta}$, trained on a source topology distribution $\mathcal{D}_{\text{train}}$, when applied directly to a target distribution $\mathcal{D}_{\text{test}}$ (where $\mathcal{D}_{\text{train}}\neq\mathcal{D}_{\text{test}}$) without parameter updates. In our experiments, this is observed by evaluating the degradation or retention of performance metrics (specifically RAC, LRC, and LAR) when the agent is transferred to unseen network topologies with varying scales and structures.
  • Scalability ($\mathcal{Sc}$): We define scalability as the algorithm’s capability to maintain efficient decision-making latency as the problem size increases. Formally, we analyze the growth rate of the inference time $T$ relative to the network size $|V|$ (number of nodes):
    $\mathcal{Sc}\propto\frac{\partial T}{\partial|V|}$ (15)
    A scalable algorithm exhibits linear or near-linear growth in $T$ rather than exponential growth, ensuring feasibility in large-scale, real-world systems with high-frequency request arrivals.
C.7 Benchmark Documentation
To ensure usability and facilitate community adoption, we provide comprehensive and well-structured documentation website for Virne project. This documentation includes detailed guides on installation, simulation configuration, API usage, and tutorials for developing new algorithms. For the anonymous requirement of the double-blind review, we include a static version of the documentation in PDF format at https://anonymous.4open.science/r/anonymous-virne/virne-document.pdf.
@@ D.1.1 Implementation Settings
Algorithm Implementation. The implementation of the RL-based NFV-RA algorithm is based on PyTorch (Paszke et al., 2019). Regarding neural policies (e.g., MLP, CNN, GNN, etc), we generally set their layers to 3 and the hidden dimension of neural networks to 128. For RL optimization, we set the reward discount factor $\lambda$ to 0.99 and use the Adam optimizer with a learning rate of 0.001 and a batch size of 128. The action section during training and testing follows a sampling strategy and a greedy strategy, respectively. Additionally, we employ the $k$-shortest paths algorithm with $k=10$ for link routing, selecting the shortest physical path that meets the bandwidth requirements for each virtual link.
Training and Testing. For RL-based methods, we train policies for each average arrival rate $\eta$, with random initialization of seeds in every simulation. The number of training epochs is set to 50, which is sufficient for all algorithms to converge. Typically, training a deep RL-based NFV-RA method completes within 5 hours. During testing, we evaluate the performance of all algorithms by repeating each test with 10 different random seeds (i.e., 0, 1111, 2222, $\ldots$, 9999) for each average arrival rate $\eta$, thereby ensuring statistical significance.
@@ D.1.2 Evaluated Physical Topologies
WX100 is a medium-scale network generated by the Waxman method (Waxman, 1988). It consists of 100 nodes connected by approximately 500 links, resulting in a density of 0.05.
@@ D.2.1 Evaluation on Varying Traffic Rates
As illustrated in Figure 10 (a), (d), and (g), there is a general downward trend in the RAC for all algorithms as the average arrival rate $\eta$ increases across all three topologies. This is an expected outcome, as higher traffic intensity leads to increased competition for finite physical network resources, inevitably resulting in more VN request rejections. As shown in Figure 10 (b), (e), and (h), many RL-based algorithms, particularly PPO-DualGAT and PPO-DualGCN, demonstrate remarkable stability in LRC across varying traffic rates. The results on LAR are depicted in Figure 10 (c), (f), and (I). Generally increasing with $\eta$, most algorithms process more VN requests. Algorithms like PPO-DualGAT and PPO-DualGCN consistently achieve higher LAR across the range of $\eta$ values, showcasing their superior capability in maximizing revenue even as conditions change. These experiments highlight that 10 while some algorithms (often sophisticated RL policies like PPO-DualGAT) exhibit graceful performance degradation and maintain high efficiency, others are more sensitive to load changes. Additionally, from the perspective of network density, it is notable that $1$ the performance difference is more significant in sparser topologies, such as BRAIN, mainly due to limited feasible action space. The denser topologies, like WX100, often have more path diversity for routing virtual links. This richer connectivity makes the embedding problem slightly less constrained, and most algorithms, including heuristics like PL-Rank, perform relatively well. However, in sparser PNs, each placement is critical because fewer alternate routes mean a poor choice can drain a key bridge’s resources and trigger subsequent failures.
Overall, this study allows researchers and practitioners to select algorithms that are well-suited to the anticipated traffic dynamics of their target environments. For instance, for networks expecting highly variable loads, algorithms demonstrating flatter RAC and LRC curves across $\eta$ would be preferable.
@@ D.2.2 Evaluation on Fluctuating Demand Distribution
As illustrated in Figure 11 (a), (d), and (g) for RAC, most algorithms exhibit an initial sharp decrease, particularly as the simulation transitions through phases with increasing VN resource demands. This phase quickly consumes available physical resources, leading to lower acceptance rates. As the simulation progresses into phases where VN sizes increase, the RAC for many algorithms continues to decline or stabilizes at a lower level. This is because larger and more complex VNs are inherently more challenging to embed. The results on R2C shown in Figure 11 (b), (e), and (h) show a dip when the demand characteristics change, but PPO-DualGCN keeps a relatively higher LRC. For WX100 ($\eta=0.04$) under changeable demands, PPO-DualGAT records an LRC of 0.703, whereas PPO-MLP has an LRC of 0.643. This suggests they are better at finding resource-efficient embeddings even when faced with unfamiliar or more challenging VN request types. The total revenue (i.e., cumulative LAR), depicted in Figure 11 (c), (f), and (i), shows a clear differentiation in the algorithms’ ability to accumulate revenue under these dynamic conditions. Algorithms that adapt well and maintain higher RAC and LRC will naturally accrue more total revenue. Notably, advanced RL methods like PPO-DualGAT and PPO-DualGCN tend to maintain a higher performance compared to others throughout these fluctuations, indicating better adaptation. From the perspective of stage evolution, in stages 1 and 2 with increased resource demands, 12 the superior performance of dual-GNN models here suggests that their learned policies latently identify action space with sufficient and rich resources within the PN to leave sufficient resources for future. Regarding stages 3 and 4 related to increased topological complexity, this mainly estimates the ability to solve a more complex graph mapping problem. The great performance of GNN-based methods further demonstrates the importance of the representation power of neural network architectures.
@@
Overall, these results highlight that 13 policies trained on a specific VN distribution may not generalize well to others, underscoring the need to generalize not only across load levels but also across demand types. For practical systems where service characteristics can evolve, choosing algorithms that demonstrate robustness in such fluctuating scenarios, as identifiable through Virne, is important.
D.3 Scalability on Network Sizes